It is found that 20% of the students failed an IELTS test. suppose 25 students are randomly selected from the total pool. Answer the following question
a) what is the probability that exactly 4 students failed the test?
b) what is the probability that more than 4 students failed the test?

Answer by stanbon(75887)

It is found that 20% of the students failed an IELTS test. suppose 25 students are randomly selected from the total pool. Answer the following question
---
Binomial Problem with n = 25 and p = 0.20
------
a) what is the probability that exactly 4 students failed the test?
P(x=4) = 25C4(0.2)^4*(0.8)^21 = 0.1867
-----------------------------------------------
b) what is the probability that more than 4 students failed the test?
P(5<= x <= 25) = 1 - P(0<= x <=4) = 1-0.4207 = 0.5793
=====================
Cheers,
Stan H.
=====================